40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu May 18 12:46:08 2017, awade, Laser, WOPO, Initial testing Innolight Diabolo SHGLenthScan_TransmissionAndPZTSig.pdf
    Reply  Sun May 21 21:11:43 2017, awade, Laser, WOPO, Initial testing Innolight Diabolo 
    Reply  Thu May 25 16:28:21 2017, Dhruva, Laser, WOPO, Initial testing Innolight Diabolo Diabolo_for_WOPO.pdf
       Reply  Fri Jun 2 15:16:00 2017, Dhruva, Laser, WOPO, Initial testing Innolight Diabolo Diabolo.pdf
Message ID: 2103     Entry time: Thu May 18 12:46:08 2017     Reply to this: 2104   2112
Author: awade 
Type: Laser 
Category: WOPO 
Subject: Initial testing Innolight Diabolo 

We are planning on using the Innolight Diabolo (given to us by Ron Drever's lab several years ago) as a source of 532 nm and 1064 nm photons for the WOPO experiment.  The status of the Diabolo is not known so I'm doing some checks.

This the 2004 edition of the Diabolo so the datasheet is difficult to find online.  I've scaned the manual + test sheets and put on the ATF wiki HERE.

---

The Diabolo unit consists of a regular 1.5 W Nd:YAG NPRO laser (pumped with two banks of diode lasers) with most of this light tapped off to an SHG unit for conversion to 532 nm.  The SHG is a hemolytic Fabry-Perot cavity made with a Lithium Niobate crystal. One end of the crystal acts as an HR/AR mirror in 1064/532 nm and the other is a HR/HR curved mirror that is mounted on a PZT.  Inside the Diabolo unit there is a EOM, MM lenses, alignment mirrors and RF+DC transmission diode to generate a PDH signal for locking.  All of the locking and cavity scanning is automatically handled by the SHG control electronics.  Details of operation are in the manual.  There is a laser control box and a separate SHG control box.  The SHG control box provides feedback control for the temperature of the crystal (nominally ~100-101 C for unpolled LiNb) and all of the modulation and demodulation electronics for PDH control. 

Initial settings on the laser, as I found it, were

Laser diode A temp = 19.69 C
Laser diode B temp = 20.39 C
Injection current  = 0 A (i.e. ramped down to off)
Laser crystal temp  = 23.42 C
 

SHG unit settings were

Double crystal temp = 102.87 C
Offset = 5.09
Gain  = 0.4
Scan amplitude  = 0 (you turn this up to scan SHG cavity)

The range of the doubling crystal temp was 102.87 C to ~ 110 C.  This seems too high for phase matching of the LN.  There is no obvious way to make a course adjustment.  Other units I've worked with (at ANU) gave a much bigger range of adjustment on SHG crystal temp centered around 100 C. We need to check the refractive indexes of LiNb for 1064 nm/532 nm as a function of temp and work out what the expect phase matching temperature is.  It might be that the temperature sensor has some kind of placement that doesn't reflect the bulk LiNb bulk temperature and the higher course offset compensates for this.  It might be worth tuning around more of the temperature space to get an idea of idea phase matching.

---

I made a few adjustments to settings on laser to bring them back to the testsheet values in the manual:
Laser diode A temp = 21.3 C
Laser diode B temp = 26.75 C
Injection current  = 2.2 A (max)
Laser crystal temp  = 23.42 C

The typical operating current was 2.4 A, but the current clamp value on the current has been reduced down to 2.2 A, this will mean a slightly lower nominal operating power.

With maximum allowable diode current (2.2 A) the output of the laser was 303 mW @ 532 nm and 302.3 mW @ 1064 nm. This was measure with a Thorlabs S302C thermal power sensor. Nominal output should be 910 mW @ 532 nm and  380 mW @ 1064 nm.  So 1064 nm is 80% of spec (expected with slightly lower current).  But the SHG 532 nm output is 33% of spec (well below).

---

As a check, I hooked up and oscilloscope to the monitor ports on the Diabolo SHG controller.  An attached plot shows a ramp of the PZT actuation signal and the 1064 nm power on transmission (measure using unit's built in PD).  Clearly there is a misalignment into the cavity, there are a bunch of HOM peaks that represent a heap of wasted light.  To fix this I need to open the Diabolo up and tweak the alignment into the SHG sub-unit.  Its not very involved but because the Eigen mode of the cavity comes to a ~25 µm waist it is highly sensitive.  Need to tweak and allow some settling time for mounts to relax. 

Attachment 1: SHGLenthScan_TransmissionAndPZTSig.pdf  104 kB  | Hide | Hide all
SHGLenthScan_TransmissionAndPZTSig.pdf
ELOG V3.1.3-